### v2

@@ Abstract
Context. Small bodies in Earth-like orbits, the Arjunas, are good targets for scientific exploration and mining studies as they enable low-cost missions. The subset of such objects that experience recurrent temporarily captured flyby or orbiter events, also called mini-moon episodes, are among the best ranked in terms of accessibility. Only a handful of objects are known to have engaged in such a dynamical behavior. Finding and characterizing more of them may help to expand scientific and commercial research activities in space over the next few decades. Asteroid 2024 PT5 was found recently and belongs to this group of interesting objects.
Aims. Here, we investigate the orbital context of 2024 PT5, and its spectral and rotational properties.
@@
Results. Asteroid 2024 PT5 experiences recurrent co-orbital engagements and mini-moon events of the temporarily captured flyby type. Its visible spectrum is consistent with that of an Sv-type asteroid or perhaps lunar ejecta. Its rotational period could be 1 h.
Conclusions. The discovery of 2024 PT5 confirms that temporarily captured flybys are relatively frequent and involve objects larger than a few meters, suitable as accessible targets for scientific research activities and commercial mining ventures in space.
@@
Both ESA and NASA are pushing into the commercial space sector supporting a nascent global space economy (see, e.g., Rausser et al. 2023; Paravano et al. 2024). Scientific exploration and mining of asteroids are activities well suited for development within the context of this fast-evolving sector (see, e.g., Sommariva 2015; Sachdeva 2018; Xie et al. 2021), but only if there are enough targets for low-cost missions. NASA’s Near-Earth Object Human Space Flight Accessible Targets Study (NHATS, Abell et al. 2012)111https://cneos.jpl.nasa.gov/nhats/ is compiling a list of near-Earth objects (NEOs) that identifies potential targets for future exploration, but such a database must include information on the composition of the targets for better mission planning. Reflectance spectroscopy can help in finding what these more accessible targets are made of (Licandro et al. 2020; de León et al. 2021).
The NHATS list includes thousands of targets, but just a few have been characterized spectroscopically (Popescu et al. 2019). Most targets with the highest number of viable trajectories (the most accessible ones) follow Earth-like paths, they are part of a secondary asteroid belt that surrounds the orbit of the Earth–Moon system. Its members define a dynamical class, the Arjunas, that follow heliocentric trajectories with orbital periods close to one sidereal year, low eccentricity, and low inclination (Rabinowitz et al. 1993). Within the 25 most accessible targets, a handful undergo recurrent temporarily captured flyby or orbiter events, also called mini-moon episodes (Granvik et al. 2012).
The recently found asteroid 2024 PT5 is part of the group of very accessible targets (de la Fuente Marcos & de la Fuente Marcos 2024). Here, we use reflectance spectroscopy, photometry, and -body simulations to confirm its nature and dynamical context. This Letter is organized as follows. In Sect. 2, we introduce the background of our research, and present the data and tools used in our analyses. In Sect. 3, we determine if 2024 PT5 is natural or artificial and its probable origin. In Sect. 4, we discuss our results and Sect. 5 summarizes our conclusions. Appendices include supporting material.
@@ 2 Context, methods, and data
In this section, we revisit dynamical concepts that are later used in our analysis. Software tools and data are also discussed here.
@@ 2.1 Dynamics background
In this work (but see Appendix A), we use the definitions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017). Temporarily captured natural irregular satellites of Earth must have negative geocentric energy and their geocentric distance must be under three Hill radii (the Hill radius of Earth is roughly 0.01 au); in other words, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1). The only NEOs that can regularly meet the capture conditions defined by Kary & Dones (1996) are the members of the Arjuna asteroid belt (Rabinowitz et al. 1993). Some of these objects experience resonant behavior induced by the 1:1 mean-motion resonance with Earth (see, e.g., de la Fuente Marcos & de la Fuente Marcos 2013). When subjected to this resonance, the relative mean longitude of the objects with respect to our planet () oscillates about a fixed value (Morais & Morbidelli 2002) and they become Earth co-orbitals as the orbital period of the NEO closely matches that of Earth. One of the resonant states compatible with co-orbital behavior drives objects to a horseshoe trajectory with respect to Earth when the orbit is plotted in a heliocentric frame of reference rotating with our planet because in this case the value of oscillates about 180°, with an amplitude 240° (Murray & Dermott 1999). In certain circumstances, perturbed horseshoe paths may lead to temporarily captured flybys as in the case of 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023) or orbiters (see Appendix B).
@@ 2.2 Data, data sources, and tools
Object A119q0V was initially reported on August 7, 2024, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Sutherland in South Africa. On August 14, it was announced with the provisional designation 2024 PT5 (Tonry et al. 2024). Its orbit determination in Table 1 is currently based on 247 observations with a data-arc span of 73 days, it is referred to epoch JD 2460600.5 TDB, which is the origin of time in the calculations, and corresponds to that of an Apollo-class asteroid. It was retrieved from Jet Propulsion Laboratory’s (JPL) Small-Body Database (SBDB)222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).333https://ssd.jpl.nasa.gov/ The object attracted our attention because it approached Earth at close range and low relative velocity.
@@
| Semimajor axis, (au) | = | 1.012305040.00000003 |
| Eccentricity, | = | 0.021476730.00000002 |
| Inclination, (°) | = | 1.5205220.000009 |
| Longitude of the ascending node, (°) | = | 305.57240.0002 |
| Argument of perihelion, (°) | = | 116.24840.0002 |
| Mean anomaly, (°) | = | 323.677200.00007 |
| Perihelion distance, (au) | = | 0.990564030.00000003 |
| Aphelion distance, (au) | = | 1.034046040.00000003 |
@@
The orbit determination of 2024 PT5 resembles that of 2022 NX1, a confirmed natural object that experiences recurrent resonant co-orbital episodes and mini-moon events of the temporarily captured flyby type. Asteroid 2022 NX1 also undergoes periodic close encounters with the Earth–Moon system that make the reconstruction of its past orbital evolution and the prediction of its future behavior beyond a few decades difficult (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023). In such cases, the orbital evolution has to be studied numerically considering the uncertainties of the orbit determination. The -body simulations needed to investigate the evolution of 2024 PT5 were carried out using a direct -body code described by Aarseth (2003) that is publicly available from the web site of the Institute of Astronomy of the University of Cambridge.555http://www.ast.cam.ac.uk/sverre/web/pages/nbody.htm This software implements the Hermite numerical integration scheme developed by Makino (1991). Technical details and relevant results from this code were discussed in de la Fuente Marcos & de la Fuente Marcos (2012). Our physical model included the perturbations by the eight major planets, the Moon, the barycenter of the Pluto-Charon system, and the 19 largest asteroids, Ceres, Pallas, Vesta, Hygiea, Euphrosyne, Interamnia, Davida, Herculina, Eunomia, Juno, Psyche, Europa, Thisbe, Iris, Egeria, Diotima, Amphitrite, Sylvia, and Doris. For accurate initial positions and velocities (see Appendix C), we used data from JPL’s SSDG Horizons on-line Solar system data and ephemeris computation service,666https://ssd.jpl.nasa.gov/horizons/ that are based on the DE440/441 planetary ephemeris (Park et al. 2021). Most input data were retrieved from JPL’s SBDB and Horizons using tools provided by the Python package Astroquery (Ginsburg et al. 2019) and its HorizonsClass class.777https://astroquery.readthedocs.io/en/latest/jplhorizons/jplhorizons.html
@@ 3 Results
Here, we use -body simulations to assess the current dynamical status of 2024 PT5, reflectance spectroscopy to determine its physical nature, and photometry to study its rotational state.
@@ 3.1 Orbital evolution
Figure 1 summarizes the results of our calculations, showing the time evolution of relevant parameters of 2024 PT5: the relative mean longitude, and the geocentric energy and distance. The figure shows the evolution of the nominal orbit and those of control orbits or clones with state vectors (Cartesian coordinates and velocities, see Table 2 in Appendix C) well away from the nominal one, up to from the nominal orbit determination in Table 1. The right panels in Fig. 1 show that the short-term evolution of all the control orbits matches that of the nominal orbit. Therefore, we confirm the results in de la Fuente Marcos & de la Fuente Marcos (2024) obtained using an early orbit determination. However, the left panel in Fig. 1 (see also Fig. 10) shows that recovering the past orbital evolution of 2024 PT5 prior to 1937 is difficult because a close encounter with Earth drives originally close orbits to diverge. A similar situation is found when predicting the dynamical behavior of this object beyond 2085. In other words, the orbital evolution over the time interval (87, 61) yr on the left panel of Fig. 1 can be computed precisely but beyond this time interval the current orbit determination cannot provide reliable ephemerides.
Figure 1, right panels, show that from 2023 until 2025, 2024 PT5 followed a horseshoe-like trajectory (top panel) not too different from that of 2022 NX1 in fig. 3 of de la Fuente Marcos et al. (2023). It is also observed that in 2024, between September 29 and November 25, the geocentric energy of all the control orbits had a negative value (Fig. 1, middle-right panel), and that its geocentric distance remained under 0.03 au during the same time interval (bottom panel). According to Kary & Dones (1996), 2024 PT5 becomes a temporary satellite of Earth during that period. Furthermore, as the asteroid fails to complete one revolution around Earth and following Granvik et al. (2012), 2024 PT5 undergoes a temporarily captured flyby. This result is statistically robust and consistent across the control orbits tested (104). Figure 1, left panels, show that 2024 PT5 might have experienced 1:1 resonant behavior of the horseshoe type in the past and perhaps it will show it again in the future (top panel), although the evolution is seriously affected by orbital chaos induced by close encounters with the Earth–Moon system. The middle and bottom panels show that recurrent temporarily captured flyby episodes could be possible. In fact, this object may experience further temporarily captured flybys in the future.
@@ 3.2 Spectroscopy
Three visible spectra of 2024 PT5 were obtained on the night of September 7, 2024, starting at 22:45 UTC, and using the OSIRIS camera spectrograph (Cepa et al. 2000; Cepa 2010). The instrument is installed at the 10.4 m Gran Telescopio Canarias (GTC), located at the El Roque de los Muchachos Observatory, in La Palma. It has a blue-sensitive 40964096 pixel CCD that yields a 7.8′7.8′ field of view. Observations were done using a slit width of 1.2″ (oriented to the parallactic angle) and the R300R grism (0.48–0.92 m). We obtained three individual spectra of 900 s of exposure time each, at an airmass of 1.45. Two solar analog stars (Landolt SA 112-1333 and SA115-271) were also observed with the same instrumental configuration and at a similar airmass to that of the asteroid, in order to get its reflectance spectrum. Both asteroid and solar analogs 2D spectral images were bias and flat-field corrected, background subtracted and collapsed to 1D, using an aperture equal to the distance from the center of the spatial profile to the pixel with 10% peak intensity. One-dimensional spectra were then wavelength calibrated with Xe+Ne+HgAr arc lamps, and a final, averaged spectrum for the asteroid and for each star were obtained. As a last step, we divided the spectrum of the asteroid by the spectrum of each solar analog star, and averaged the two resulting ratios. The final reflectance spectrum of 2024 PT5 is shown in Fig. 2, in green.
@@
We used the M4AST888https://spectre.imcce.fr/m4ast/index.php/index/home online tool (Popescu et al. 2012) to get the most representative taxonomical type for 2024 PT5. This tool fits a curve to the spectrum and compares it to the ‘taxons’ defined by DeMeo et al. 2009 using a procedure. The best fit corresponds to an Sv-type (in orange in Fig. 2), a class that serves as a link between S-types (mixtures of silicates and metal) and V-types, indicative of basaltic surfaces, and with asteroid (4) Vesta as the main representative of the class. The M4AST tool was also used to compare the spectrum of the asteroid with more than 15 000 spectra of meteorites, terrestrial rocks, and lunar soils contained in the RELAB999https://sites.brown.edu/relab/relab-spectral-database/ database. Figure 2 shows the best fit in red, corresponding to a powder sample (grain sizes from 10 to 45 m) from a mare breccia of the Moon, collected by the Luna 24 mission. Interestingly, and in order of increasing values of , this best fit is followed by several samples from Moon maria collected at the Apollo 12 landing site, and a sample of a terrestrial basaltic rock (Brockenheim, Germany).
@@
Sharkey et al. 2021 found that Earth’s quasi-satellite (469219) 2016 HO3 Kamo’oalewa (de la Fuente Marcos & de la Fuente Marcos 2016) presents a visible spectrum consistent with S- and L-type taxonomic classes, but when extended to the near-infrared, it shows a very steep spectral slope, consistent with space-weathered, lunar-like silicates. The visible spectrum of Kamo’oalewa (in light-blue), shown in Fig. 3 and digitized from Sharkey et al. 2021, is similar to that of 2022 NX1 as published in de la Fuente Marcos et al. (2023) and classified as a K-type, although we would need the near-infrared to confirm such an agreement. Note that the visible spectrum of 2024 PT5 shows a deeper absorption band at 1 m, compared to those from Kamo’oalewa and 2022 NX1, suggesting slightly less weathered silicates for the former.

### v3

@@ Abstract
Context. Small bodies in Earth-like orbits, the Arjunas, are good targets for scientific exploration, and space mining or in situ resource utilization (ISRU) trials as they enable low-cost missions. The subset of these objects that experience recurrent temporarily captured flyby or orbiter episodes are among the best ranked in terms of accessibility. Only a handful of objects are known to have engaged in such a dynamical behavior. Finding and characterizing more of them may help to expand scientific and commercial research activities in space over the next few decades. Asteroid 2024 PT5 is a recent finding that shows dynamical traits in common with this group.
Aims. Here we investigate the orbital context of 2024 PT5 and its spectral and rotational properties.
@@
Results. Asteroid 2024 PT5 experiences recurrent co-orbital engagements and episodes in which it has negative geocentric orbital energy while inside a geocentric distance under three Hill radii, which we call quasi-mini-moon events. Its visible spectrum is consistent with that of lunar-like silicates. Photometric data suggest a rotation period 1 h.
Conclusions. The discovery of 2024 PT5 confirms that events resembling temporary captures are relatively frequent and involve objects larger than a few meters, suitable as accessible targets for scientific research activities and demonstrating ISRU technologies.
@@
Both ESA and NASA are pushing into the commercial space sector supporting a nascent global space economy (see, e.g., Rausser et al. 2023; Paravano et al. 2024). Scientific exploration and mining or in situ resource utilization (ISRU) of asteroids are activities well suited for development within the context of this fast-evolving sector (see, e.g., Granvik et al. 2013; Sercel et al. 2018; Xie et al. 2021), but only if there are enough targets for low-cost missions. NASA’s Near-Earth Object Human Space Flight Accessible Targets Study (NHATS; Abell et al. 2012)111https://cneos.jpl.nasa.gov/nhats/ is compiling a list of near-Earth objects (NEOs) that identifies potential targets for future exploration, but such a database must include information on the composition of the targets for better mission planning. Reflectance spectroscopy can help in finding what these more accessible targets are made of (Licandro et al. 2020; de León et al. 2021).
NHATS includes thousands of targets, but few have been characterized spectroscopically (Popescu et al. 2019). Most targets with the highest number of viable trajectories (the most accessible ones) follow heliocentric orbits with periods close to one sidereal year and have low eccentricity and low inclination. They are part of a secondary asteroid belt that surrounds the orbit of the Earth–Moon system and define a dynamical class, the Arjunas (Rabinowitz et al. 1993). Within the 25 most accessible targets, a handful have recurring temporarily captured flyby or orbiter events, also called mini-moon episodes (Granvik et al. 2012).
The recently found asteroid 2024 PT5 is part of the group of very accessible targets (de la Fuente Marcos & de la Fuente Marcos 2024). Here we used reflectance spectroscopy, photometry, and -body simulations to confirm its nature and dynamical context. This Letter is organized as follows. In Sect. 2 we introduce the background of our research, and present the data and tools used in our analyses. In Sect. 3 we investigate whether 2024 PT5 is natural or artificial and its probable origin. In Sect. 4 we discuss our results. Section 5 summarizes our conclusions. The Appendices include supporting material.
@@ 2 Context, methods, and data
In this section we revisit the dynamical concepts that are used in our analysis. Software tools and data are also discussed here.
@@ 2.1 Dynamics background
For this work (but see Appendix A) we adopted the criteria in Kary & Dones (1996) and Granvik et al. (2012). Temporarily captured natural irregular satellites of Earth must have negative geocentric energy, and their geocentric distance must be under three Hill radii (Earth’s Hill radius is 0.01 au); in other words, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1). The only NEOs that can regularly meet the capture conditions defined by Kary & Dones (1996) are the Arjunas. Some of them also experience resonant behavior induced by the 1:1 mean-motion resonance (see, e.g., de la Fuente Marcos & de la Fuente Marcos 2013) in which their relative mean longitude with respect to our planet () oscillates about a fixed value (Morais & Morbidelli 2002) to become Earth co-orbitals as their orbital periods closely match that of Earth. One of the resonant states compatible with co-orbital behavior drives objects to a horseshoe trajectory when plotted in a heliocentric frame of reference rotating with our planet because in this case the value of oscillates about 180°, with an amplitude 240° (Murray & Dermott 1999). Perturbed horseshoe paths may lead to temporarily captured flybys, as in the case of 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023), or orbiters (see Appendix B).
@@ 2.2 Data, data sources, and tools
Object A119q0V was initially reported on August 7, 2024, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Sutherland in South Africa. It was announced on August 14 with the provisional designation 2024 PT5 (Tonry et al. 2024). Its orbital solution in Table 1 is consistent with that of an Apollo-class asteroid. It is currently based on 358 observations with an observational timespan of 172 days and referred to epoch JD 2460600.5 TDB, which is the origin of time in the calculations. It was retrieved from the Jet Propulsion Laboratory (JPL) Small-Body Database (SBDB)222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).333https://ssd.jpl.nasa.gov/ The object attracted our attention because it approached Earth at close range and low relative velocity.
@@
| Semimajor axis, (au) | = | 1.0123050600.000000009 |
| Eccentricity, | = | 0.0214767030.000000006 |
| Inclination, (°) | = | 1.52051670.0000004 |
| Longitude of the ascending node, (°) | = | 305.5723600.000007 |
| Argument of perihelion, (°) | = | 116.248440.00003 |
| Mean anomaly, (°) | = | 323.677250.00003 |
| Perihelion distance, (au) | = | 0.9905640840.000000005 |
| Aphelion distance, (au) | = | 1.0340460350.000000009 |
@@
The orbit of 2024 PT5 resembles that of 2022 NX1, a confirmed natural object that experiences recurrent resonant co-orbital episodes and temporarily captured flybys. Asteroid 2022 NX1 also undergoes periodic close encounters with the Earth–Moon system that make the reconstruction of its past orbital evolution and the prediction of its future behavior beyond a few decades difficult (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023). In such cases, the orbital evolution has to be studied statistically considering the uncertainties of the orbit. The calculations needed to investigate the evolution of 2024 PT5 were carried out using a direct -body code described by Aarseth (2003), publicly available from the website of the Institute of Astronomy of the University of Cambridge,555https://people.ast.cam.ac.uk/ sverre/web/pages/nbody.htm which implements the Hermite numerical integration scheme developed by Makino (1991). Further details and the relevant results from this code were discussed in de la Fuente Marcos & de la Fuente Marcos (2012). Our physical model included the perturbations by the eight major planets, the Moon, the barycenter of the Pluto-Charon system, and the 19 largest asteroids: Ceres, Pallas, Vesta, Hygiea, Euphrosyne, Interamnia, Davida, Herculina, Eunomia, Juno, Psyche, Europa, Thisbe, Iris, Egeria, Diotima, Amphitrite, Sylvia, and Doris. For accurate initial positions and velocities (see Appendix C), we used data based on the DE440/441 planetary ephemeris (Park et al. 2021) from JPL’s SSDG Horizons online Solar System data and ephemeris computation service.666https://ssd.jpl.nasa.gov/horizons/ Most input data were retrieved from SBDB and Horizons using tools provided by the Python package Astroquery (Ginsburg et al. 2019) and its HorizonsClass class.777https://astroquery.readthedocs.io/en/latest/jplhorizons/jplhorizons.html
@@ 3 Results
We used -body simulations to assess the current dynamical status of 2024 PT5, reflectance spectroscopy to determine its physical nature, and photometry to study its rotational state.
@@ 3.1 Orbital evolution
Figure 1 summarizes the results of our calculations, showing the time evolution of the relevant parameters of 2024 PT5: the relative mean longitude, and the geocentric energy and distance. The figure shows the evolution of the nominal orbit and those of the control orbits or clones with state vectors (Cartesian coordinates and velocities; see Table 2) well away from the nominal orbit, up to from the nominal orbital solution in Table 1. The right panels in Fig. 1 show that the short-term evolution of all the control orbits matches that of the nominal orbit. Therefore, we confirm the results in de la Fuente Marcos & de la Fuente Marcos (2024) obtained using an early orbital solution. However, the left panel in Fig. 1 (see also Fig. 10) shows that recovering the past orbital evolution of 2024 PT5 prior to 1937 is difficult because a close encounter with Earth causes originally close orbits to diverge. This is also found when predicting the behavior of this object beyond 2084. In other words, the orbital evolution over the time interval (87, 60) yr in the left panel of Fig. 1 can be computed precisely, but beyond this time interval the current orbital solution cannot provide reliable ephemerides.
Figure 1, right panels, shows that from 2023 until 2025, 2024 PT5 followed a horseshoe-like trajectory (top panel) not too different from that of 2022 NX1 in Fig. 3 of de la Fuente Marcos et al. (2023). It is also observed in 2024, between September 29 and November 25, that the geocentric energy of all the control orbits had a negative value (Fig. 1, middle right panel), and that its geocentric distance remained under 0.03 au during the same time interval (bottom panel). According to Kary & Dones (1996), 2024 PT5 was a temporary satellite of Earth during that period. Furthermore, as the asteroid failed to complete one revolution around Earth and following Granvik et al. (2012), 2024 PT5 underwent a temporarily captured flyby; we call this a quasi-mini-moon event. This result is statistically robust and consistent across the control orbits tested (104). Figure 1, left panels, shows that the evolution of 2024 PT5 is seriously affected by orbital chaos induced by close encounters with the Earth–Moon system. The middle and bottom panels show that recurrent temporarily captured flyby episodes could be possible. In fact, this object may experience further temporarily captured flybys in the future (in 2084 for about 43 d).
@@ 3.2 Spectroscopy
Three visible spectra of 2024 PT5 were obtained on the night of September 7, 2024, starting at 22:45 UTC, using the OSIRIS camera spectrograph (Cepa et al. 2000; Cepa 2010). The instrument is installed at the 10.4 m Gran Telescopio Canarias (GTC), located at the El Roque de los Muchachos Observatory, on La Palma. It has a blue-sensitive 40964096 pixel CCD that yields a 7.8′7.8′ field of view. Observations were done using a slit width of 1.2″ (oriented to the parallactic angle) and the R300R grism (0.48–0.92 m). We obtained three individual spectra of 900 s of exposure time each, at an airmass of 1.45. Two solar analog stars (Landolt SA 112-1333 and SA115-271) were also observed with the same instrumental configuration and at a similar airmass to that of the asteroid in order to get its reflectance spectrum. The 2D spectral images of the asteroid and the two solar analogs were bias and flat-field corrected, background subtracted, and collapsed to 1D using an aperture equal to the distance from the center of the spatial profile to the pixel with 10% peak intensity. One-dimensional spectra were then wavelength calibrated with Xe+Ne+HgAr arc lamps, and a final averaged spectrum for the asteroid and for each star were obtained. As a last step, we divided the spectrum of the asteroid by the spectrum of each solar analog star, and averaged the two resulting ratios. The final reflectance spectrum of 2024 PT5 is shown in Fig. 2, in green.
@@
We used the M4AST888https://spectre.imcce.fr/m4ast/index.php/index/home online tool (Popescu et al. 2012) to get the most representative taxonomical type for 2024 PT5. This tool fits a curve to the spectrum and compares it to the “taxons” defined by DeMeo et al. (2009) using a procedure. The best fit corresponds to an Sv-type (in orange in Fig. 2), a class that serves as a link between S-types (mixtures of silicates and metal) and V-types, indicative of basaltic surfaces, and with asteroid (4) Vesta as the main representative of the class. The M4AST tool was also used to compare the spectrum of the asteroid with more than 15 000 spectra of meteorites, terrestrial rocks, and lunar soils contained in the RELAB999https://sites.brown.edu/relab/relab-spectral-database/ database. Figure 2 shows the best fit in red, corresponding to a powder sample (grain sizes from 10 to 45 m) from a mare breccia of the Moon, collected by the Luna 24 mission. Interestingly, and in order of increasing values of , this best fit is followed by several samples from Moon maria collected at the Apollo 12 landing site, and a sample of a terrestrial basaltic rock (Brockenheim, Germany).
@@
Sharkey et al. (2021) found that Earth’s quasi-satellite (469219) 2016 HO3 Kamo’oalewa (de la Fuente Marcos & de la Fuente Marcos 2016) has a visible spectrum consistent with the S- and L-type taxonomic classes (see Fig. 3), similar to that of 2022 NX1 published in de la Fuente Marcos et al. (2023) and classified as a K-type, and different from the visible spectrum of 2024 PT5. However, when extended to the near-infrared, Kamo’oalewa shows a very red spectral slope, consistent with space-weathered lunar-like silicates. This outlines the importance of having spectral information at near-infrared wavelengths, even in the form of colors, as in Sharkey et al. (2021), to better constrain the surface composition of asteroids. We note that our visible spectrum of 2024 PT5 shows a deeper absorption band at 1 m compared to that of Kamo’oalewa, suggesting less weathered lunar-like silicates.